OpenStudy (mwhit2101):

Sixth Grade Probability (Independent and Dependent Events) Help?? Medal+Fan! A bag contains 5 chocolate, 3 peanut butter, 4 oatmeal and 4 sugar cookies. If 2 cookies are selected randomly, what is the probability that they will be the same kind?

OpenStudy (anonymous):

You already know there are 16 cookies, now you have to set up the problem and solve it

OpenStudy (anonymous):

P(both chosen are peanut butter cookies), Now with that provided you should be able to write out the probability for two of the same cookies being picked, tell me if you need help

OpenStudy (mwhit2101):

I don't get it, because wouldn't the probability change between the numbers of the cookies? Like, there's 5 chocolate but only 3 peanut butter, etc... or I guess not because it's two for everything? But I still don't get how to get the probability @Night-Watcher

OpenStudy (anonymous):

3/16 * 2/15 Multiply them two, then simplify the answer.

OpenStudy (mwhit2101):

1/40 @Night-Watcher

OpenStudy (mwhit2101):

Ok...but that doesn't answer what the probability is? 1/40 is none of my choices.

OpenStudy (mwhit2101):

My choices are: a. 5/24 b. 5/7 c. 5/12

OpenStudy (mwhit2101):

Someone heeeeelp :(

OpenStudy (anonymous):

u should use ppermutations and combinations example: selecting 2 from peanut is 3C2/16C2 Similarly 2 of any kind is 3C2+5C2+4C2+4C2.it should be divided by 16C2

OpenStudy (mwhit2101):

What do I have to do then? Divide all of those?

OpenStudy (anonymous):

nCr=n!/(r!*(n-r)I so 3C2=3 5C2=20 4C2=6 add above and divide it by 16C2 which is 120

OpenStudy (anonymous):

sorry 5C2=10

OpenStudy (anonymous):

are u able to understand???

OpenStudy (anonymous):

(3+10+6+6)/120

OpenStudy (mwhit2101):

(15)/120?

OpenStudy (mwhit2101):

8.

OpenStudy (anonymous):

25/120

OpenStudy (mwhit2101):

thats what i meant oops

OpenStudy (mwhit2101):

4 with 8 remaining

OpenStudy (anonymous):

25/120=5/24

OpenStudy (mwhit2101):

agh sorry i keep typing in wrong number on my calculator

OpenStudy (mwhit2101):

Ok 5/24 is the answer? Thanks! Can you explain to me what all that C stuff was all about so I leave with something because I didn't really get what you did

OpenStudy (anonymous):

actually, nCr is selecting r objects in total of "n". SO if u want to select 2 obj out of 10 then it is 10C2. nCr formula=n!/(r!*(n-r)!) so 10C2=10!/(2!*8!) where 10!=10*9*8*....2*1 Here in this sum u have to select 2 of same kind which can be either peanut,chocolate,cookies or oatmeal. hence it is 5C2+3C2+4C2+4C2. total bag have 16 cookies.SO selecting any two is 16C2 Answer is( 5C2+3C2+4C2+4C2)/16C2

OpenStudy (anonymous):

if u still did not understand ask me i wil help u